OpenStudy (anonymous):

You are standing on a scale inside an elevator. The scale reads 59 percent of your weight. What is the acceleration of the elevator?(Define positive y upward explain how to do it please

OpenStudy (anonymous):

OK first, what direction does the elevator need to be going for you to be lighter?

OpenStudy (anonymous):

^^^^ down

OpenStudy (anonymous):

Answer 5.79 m/s^2 if using 9.81 m/s^2 for gravity 5.9 m/s^2 if using 10m/s^2 for gravity.

OpenStudy (anonymous):

How to solve: Give a fictional weight to your situation, lets say you weigh 50kg, now what weight would the scale normal say in Newtons? F=mg=(10)(50)=500N (using 10 for gravity) now multiply that by the % of weight it reads, which is 59%, so .59 500 x (.59) = 295N therefore F=ma=295N (50kg)(a) = 295N a = 295/50 a = 5.9 m/s^2